OpenStudy (mathmath333):

partial fraction \(\large\tt \color{black}{\dfrac{x^5-1}{x^2(x^3+1)}}\)

OpenStudy (anonymous):

\[\frac{x^5-1}{x^2 \left(x^3+1\right)}=\frac{2 x-1}{3 \left(x^2-x+1\right)}-\frac{1}{x^2}-\frac{2}{3 (x+1)}+1 \]

OpenStudy (mathmath333):

i dont get this can u tell the method

OpenStudy (mathmath333):

@iambatman ,

ganeshie8 (ganeshie8):

first step is to make numerator's degree lesser than denominator's

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{x^5-1}{x^2(x^3+1)}=1-\dfrac{x^2}{x^2(x^3+1)}}\)

ganeshie8 (ganeshie8):

thats a good start, but nobody would leave x^2 like that without canceling

OpenStudy (mathmath333):

this one \(\large\tt \color{black}{\dfrac{x^5-1}{x^2(x^3+1)}=1-\dfrac{(x^2+1)}{x^2(x^3+1)}}\)

ganeshie8 (ganeshie8):

looks good

ganeshie8 (ganeshie8):

\[\large \dfrac{(x^2+1)}{x^2(x^3+1)} =\dfrac{(x^2+1)}{x^2(x+1)(x^2-x+1)} \]

ganeshie8 (ganeshie8):

just factored x^3+1 using a^3+b^3 formula ^^

ganeshie8 (ganeshie8):

see if that makes sense so far

OpenStudy (mathmath333):

yes i got till now

ganeshie8 (ganeshie8):

good, next refer this page http://www.mathsisfun.com/algebra/partial-fractions.html

ganeshie8 (ganeshie8):

\[\large \dfrac{(x^2+1)}{x^2(x+1)(x^2-x+1)} = \dfrac{A}{x} + \dfrac{B}{x^2} + \dfrac{C}{x+1} + \dfrac{Dx+E}{x^2-x+1} \]

ganeshie8 (ganeshie8):

solve the constants A,B,C,D,E

OpenStudy (mathmath333):

in place of B should be Bx + constant here ?

OpenStudy (mathmath333):

because x^2 ?

OpenStudy (anonymous):

Regarding the \(\dfrac{\cdots}{x^2}\) term... If you set the numerator to a linear factor \(bx+c\), you'll eventually find that \(b=0\), so it's the same as simply using the constant factor \(B\) as @ganeshie8 has it. For future reference, given a factor with multiplicity > 1, you can use the same numerator as for lower order terms. Examples: (Let \(P(x)\) be a polynomial) \[\frac{P(x)}{x^2}=\frac{A}{x}+\frac{B}{x^2}\] \[\frac{P(x)}{(x+1)^2}=\frac{A}{x+1}+\frac{B}{(x+1)^2}\] \[\frac{P(x)}{x^2(x^2+1)^2}=\frac{A}{x}+\frac{B}{x^2}+\frac{Cx+D}{x^2+1}+\frac{Ex+F}{(x^2+1)^2}\]

OpenStudy (mathmath333):

that i was missing thanks very much both